OpenStudy (anonymous):

√x-1 -x= -1 ( what is x??)

OpenStudy (anonymous):

\[\sqrt{x-1}= x-1\]R u there? I can show you how to do this

OpenStudy (anonymous):

yes!!! im here!!!

OpenStudy (anonymous):

Ok so first you have to get the radical by itself which i did by adding x to each side of teh equation

OpenStudy (anonymous):

Now we have to square both sides to get rid of the radical

OpenStudy (anonymous):

\[x-1=(x-1)(x-1)\]Careful because you have to FOIL the right side

OpenStudy (anonymous):

\[x-1=x^2-2x+1\]

OpenStudy (anonymous):

Now set the equation equal to zero.

OpenStudy (anonymous):

sure.... yes!!!

OpenStudy (anonymous):

X^2-x +2

OpenStudy (anonymous):

\[0=x^2-3x+2\]You can either factor this and solve for x or use the quadratic

OpenStudy (anonymous):

Right??? In algebra 2, were learning how to factor radicals....teach me how to do that!!!

OpenStudy (anonymous):

\[0=(x-2)(x-1)\]\[x=2,1\]Here's the last step. Now I'll teach you how to do it

OpenStudy (anonymous):

k

OpenStudy (anonymous):

 Right??? Thank you!!! You've been a great help.

OpenStudy (anonymous):

\[x^2-3x+2\]So lets factor this. You want to break up the middle term into 2 numbers that add to the middle term(-2-1=-3) Also these numbers when multiplied should equal the last term(-2*-1=2).\[x^2-2x-x+2\]Now we are going to group them\[(x^2-2x)+(-x+2)\]Now we have oto do some small factoring\[x(x-2)+-1(x-2)\]x is in each term so I factored that out. I wanted a positive x so i factored out -1 from the other term. Now we have a (x-2) in each term so we are actually going to factor this out.\[(x-2)(x-1)\]Done. Questions?

OpenStudy (anonymous):

correct to ur previous comment

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Did you understand the factoring? Eventually you'll get good and recognize patterns and not have to do any of that